2.A local landscape company had a 1-week sale.
On Monday,Mrs.Jones had $82 to spend.After purchasing 5 trees,she had just enough money left to purchase 1 shrub.Later in that same week,she purchased 2 trees.She had $37 with her,so she again had enough money left to purchase 1 shrub.Find the cost of a tree and the cost of a shrub.

Answer by stanbon(75887)

On Monday,Mrs.Jones had $82 to spend.After purchasing 5 trees,she had just enough money left to purchase 1 shrub.Later in that same week,she purchased 2 trees.She had $37 with her,so she again had enough money left to purchase 1 shrub.Find the cost of a tree and the cost of a shrub.
------------------------------------------
EQUATIONS:
5t+s = 82
2t+s = 37
------------
Subtract the 2nd equation from the 1st to get:
3t = 45
t = $15 (price of each tree)
-------
Substitute to solve for "s":
2*15 + s = 37
s = $7 (price of each shrub)
===================
Cheers,
Stan H.